OpenStudy (anonymous):

For exercises 4-5 use this equation to estimate typing speed S= W/5 – 2e, where S is the accurate typing speed, W is the number of words typed in 5 minutes, and E is the number of errors. Tanner usually makes 3 errors every 5 minutes when he is typing. If his goal is an accurate typing speed of at least 55 words per minute, how many words does he have to be able to type in 5 minutes? A.) 325 words B.) 300 words C.) 305 words D.) 61 words

OpenStudy (anonymous):

D: another one!

jimthompson5910 (jim_thompson5910):

S= W/5 – 2e 55= x/5 – 2*3 55= x/5 – 6 55+6 = x/5 61 = x/5 I'll let you finish

OpenStudy (anonymous):

305?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks(:

jimthompson5910 (jim_thompson5910):

np